Show that z = i is a root of the equation z4 +z3 +z−1 = 0. Find the three other roots.
1
Expert's answer
2020-05-24T16:03:19-0400

Equation z4 +z3 +z−1 = 0 can be written after re-arranging the terms

=> (z4-1) +z(z2+1)=0

=> (z2-1)(z2+1)+z(z2+1)=0

=> (z2+1) (z2+z-1) =0

=> So roots of z2 +1 =0 are -i and i;

roots of (z2 +z-1) =0 are "(-1+\sqrt{\smash[b]{5}})/2" and "(-1-\sqrt{\smash[b]{5}})/2".

hence other 3 roots are  -i, "(-1+\sqrt{\smash[b]{5}})/2" and "(-1-\sqrt{\smash[b]{5}})/2" .